# Exploring nuclear force with pulsar glitch observation

### v1

@@ Abstract
We connect nuclear forces to one of the most notable irregular behaviors observed in pulsars, already detected in approximately 6% known pulsars, with increasingly accurate data expected from upcoming high-precision timing instruments on both ground and space. Built on Shang & Li (2021), we conduct a case study on the 2001 glitch of the Vela pulsar. For our purpose, we adopt the Relativistic Mean Field (RMF) model as the theoretical many-body framework to describe nuclear systems. We refit three representative RMF parameter sets (DD-ME2, PKDD, NL3), considering the uncertainties in nuclear matter saturation properties. Utilizing the resulting star structure, composition and nucleon properties in the medium obtained in a consistent manner, we calculate the pinning energy of superfluid vortex in the nuclear lattice in the inner crust. This leads to the evolution of associated pinning force that acts on the vortex, which can be confronted with observed glitch amplitude and short-time relaxation in the 2000 Vela glitch event, following the superfluidity model of pulsar glitch. We discuss how the vortex configuration and pinning properties depend on the nuclear parameters, and find an interesting and dominant role of the nuclear symmetry energy slope on pinning strength.
@@ 1 Introduction
This paper is organized as follows. In Sec. 2.12.5, the theoretical framework for constructing unified EoS, the composition of NSs, the neutron pairing gap, and the pinning energy are given, respectively. Sec. 2.6 is dedicated to formulating the snowplow model. In Sec. 3, we demonstrate the dependence of pinning energy and pinning force on the effective interactions and polarization, and present the constraint on the nuclear parameters and Vela mass with the Vela 2000 glitch observation. Some necessary discussions are also in Sec. 3. Finally, a brief summary is given in Sec. 4.
@@ 2.1 RMF framework for nuclear matter and refitting parameter sets
For describing nuclear matter, we consider the baryons interacting with each other through the exchange of isoscalar scalar and vector mesons ($\sigma$ and $\omega$), isovector vector meson ($\rho$) in the RMF model. The Lagrangian density that describes the systems with time-reversal symmetry can be written as: {widetext}
$\mathcal{L}=$ $\sum_{N=n,p}\bar{\psi}_{N}\left\{\gamma^{\mu}\left[i\partial_{\mu}-g_{\omega N}\omega_{\mu}-g_{\rho N}\boldsymbol{\rho}_{\mu}\boldsymbol{\tau}_{N}-q_{i}A_{\mu}\right]-\left[M_{N}-g_{\sigma N}\sigma\right]\right\}\psi_{N}$ (1)
@@
where $\boldsymbol{\tau}_{N}$ and $q_{i}$ are the Pauli matrices and charges of nucleons, and $M_{N}$ and $m_{l}$ represent the nucleon and lepton masses, respectively. $g_{mN}$ is the coupling constant between the meson $m$ and the nucleon $N$. $\psi_{N(l)}$ is the Dirac field of the nucleons or the leptons. $\sigma$, $\omega_{\mu}$, and $\boldsymbol{\rho}_{\mu}$ denote the quantum fields of mesons. The field tensors of $\omega$, $\rho$, and photon are
@@
The interaction term in the RMF Lagrangian Eq. (1) depends on the nucleon-meson coupling constants that are usually determined by fitting nuclei or nuclear-matter properties. For a quantitative description of nuclear matter and finite nuclei, one needs to include a medium dependence of the effective mean-field interactions accounting for higher-order many-body effects. The medium dependence can either be introduced by including non-linear (NL) meson self-interaction terms in the Lagrangian, or by assuming an explicit density dependence (DD) for the meson-nucleon couplings. In the NLRMF model, self-coupling terms and cross-coupling terms for the mesons are introduced in $U(\sigma,\omega_{\mu},\boldsymbol{\rho}_{\mu})$ with the flexibility of adjusting the density dependence of the symmetry energy, defined based on the parabolic approximation of the binding energy per nucleon $E/A=\varepsilon/\rho-M_{N}$: $E/A(\rho_{\rm B},\beta)=E/A(\rho_{\rm B},\beta=0)+E_{\rm sym}(\rho_{\rm B})\beta^{2}+...$ with $\beta=(\rho_{n}-\rho_{p})/\rho_{\rm B}$ denoting the isospin asymmetry. Density-dependent coupling parameters in the DDRMF model correspond to $U(\sigma,\omega_{\mu},\boldsymbol{\rho}_{\mu})$ in the NLRMF model. The density dependence of the coupling can be in phenomenological functional form, with parameters adjusted to experimental data. For example, following the Typel-Wolter ansatz, the density dependence of the coupling parameters are
@@
The RMF model provides an excellent tool to study the properties of finite nuclei and infinite nuclear matter (Walecka, 1974; Glendenning, 1996; Shen, 2002; Dutra et al., 2014; Li et al., 2020; Rong et al., 2021; Tu & Zhou, 2022; Sun et al., 2023; Zhu et al., 2023). See Dutra et al. (2014) for a recent comparison of different RMF parameterizations. The extensions of several effective interactions, e.g., DD-ME2 (Lalazissis et al., 2005; Li & Sedrakian, 2019), PKDD (Long et al., 2004), and NL3 (Schaffner-Bielich & Gal, 2000; Wu et al., 2021), can be utilized to study the impact of nuclear symmetry energy on the pinning properties, with different density-slope of the symmetry energy at saturation, $L_{0}=3\rho_{0}(\frac{\mathrm{d}E_{\rm sym}}{\mathrm{d}\rho_{\rm B}})|_{\rho_{0}}$.
Table. 1 lists the characteristic coefficients of nuclear matter for the original effective interactions DD-ME2, PKDD, and NL3. The original DD-ME2 is a widely-adopted interaction that effectively describes the properties of finite nuclei and nuclear matter, having the smallest symmetry energy slope $L_{0}=$ 51.26 MeV among the three effective interactions. In contrast, the original NL3 is characterized by an excessively large symmetry energy slope of $L_{0}=$ 118.5 MeV.
@@
Table. 2 presents the coupling parameters $g_{\rho}(\rho_{0})$ and $a_{\rho}$ for different effective interactions with $L_{0}=$30, 40, 60, and 80 MeV. NL3 with $L_{0}=$ 30 MeV is excluded from our work due to its non-monotonic behavior in the EoS.
@@ 2.2 Unified description of stellar crust and core
For the uniform nuclear matter in the core, due to the translational invariance of the system, all derivative terms in Eqs. (3)–(6) vanish. Using the Thomas-Fermi approximation, the nucleon Dirac field can be considered as plane wave, leading to a nucleon single-particle energy,
@@
The value of $E_{N}(k)$ at the Fermi surface is the corresponding chemical potential. $\Sigma_{R}$ is the rearrangement term that originated from the density dependence of the coupling constants between nucleons and mesons, $\Sigma_{R}=0$ for the NLRMF model. By combining the $\beta$-equilibrium condition, the charge neutrality condition, and the nucleon number conservation condition in the core, the nucleon Fermi momentum and meson fields at a given number density can be self-consistently solved.
Furthermore, the energy density $\varepsilon$ and pressure $P$ can be calculated by using the energy–momentum tensor
@@
For the non-uniform matter of the outer and inner crusts, the energy density $\varepsilon_{\rm{non}}$ is calculated by integrating the whole WS cell,
@@
Apply the spherical Wigner-Seitz approximation and reflective boundary condition, the distributions of meson fields and densities of nucleons and leptons are obtained by iteratively solving the nonlinear system of equations containing Eqs. (3-6), $\beta$-equilibrium, charge neutrality, and nucleon number conservation conditions in the spherical symmetry one-dimensional geometry. In each iteration, the meson field is expanded by fast cosine transformation to improve the computational efficiency. The non-uniform structure is stable after fulfilling the constancy of chemical potentials,
@@
The vertical lines on the left and right sides of Fig. 1 represent the outer-inner crusts transition densities $\rho_{\rm{oi}}$ and crust-core transition densities $\rho_{\rm{cc}}$ derived from different effective interactions, respectively. Table. 3 lists their exact values. The outer-inner crust transition density $\rho_{\rm{oi}}$ is fixed by the chemical potential of neutron exceeding its rest mass $M_{N}$, while the crust-core transition density $\rho_{\rm{cc}}$ is determined by $\varepsilon_{\rm{uni}}$ being less than $\varepsilon_{\rm{non}}$. It is clearly that as $L_{0}$ increases, $\rho_{\rm{oi}}$ increases while $\rho_{\rm{cc}}$ decreases, consistent with previous works (Fantina et al., 2016).
@@ 2.3 Global properties and composition of neutron stars
Figure 4: (a) $M$-$R$ relations and (b) mass profiles calculated with the unified EoSs of DD-ME2 (solid lines), PKDD (dashed lines), and NL3 (dotted lines) with the different symmetry energy slope $L_{0}=$30, 40, 60, and 80 MeV. The $1\sigma$ credible regions of joint mass-radius measurements from GW observations for GW170817 and X-ray observations for PSR J0030+0451, PSR J0740+6620 and PSR J0437-4715 are shaded in panel (a). The dashed horizontal line presents $2M_{\odot}$ in panel (b).
Now we discuss the dependence of droplet size $R_{\rm{d}}$ and WS cell size $R_{\rm{WS}}$ on effective interactions, particularly $L_{0}$. Fig. 3 illustrates $R_{\rm{d}}$ and $R_{\rm{WS}}$ calculated using DD-ME2 with $L_{0}=$30, 60, and 80 MeV as a function of $\rho_{\mathrm{B}}$. At lower densities, $L_{0}$ has a minor effect on $R_{\rm{d}}$ and $R_{\rm{WS}}$. However, near the crust-core transition densities, we can observe a dramatic increase in both $R_{\rm{d}}$ and $R_{\rm{WS}}$; this effect becomes more pronounced with smaller $L_{0}$ and tends to disappear at $L_{0}=$ 80 MeV. The effects of symmetry energy on $R_{\rm{d}}$ and $R_{\rm{WS}}$ can be explained by the surface tension. Oyamatsu & Iida (2007) has showed that the surface tension decreases with increasing $L_{0}$. Consequently, droplets calculated at lower $L_{0}$ can maintain inhomogeneous structures under higher density and stronger Coulomb force due to the larger surface tension, thus further supporting a higher transition density and a larger droplet size. The upper panel of Fig. 3 for $L_{0}=$ 30 MeV suggests the possible existence of the giant droplet and extremely dense inhomogeneous structure near the crust-core transition density, which could significantly reduce pinning energy.
Both the global properties and composition of NSs are essential to simulations of pulsar glitches. Using the unified EoS shown in Fig. 1 as basic input, the global properties of NSs, e.g., $M$-$R$ relation and mass profile, can be obtained by solving the Tolman–Oppenheimer–Volkoff (TOV) equation (Tolman, 1939; Oppenheimer & Volkoff, 1939),
@@
where $r$ is the distance from the center of NSs. Given a central density $\rho_{c}$ at $r$, the TOV equation is integrated from $r=0$ to $r=R$ where the pressure is zero. $R$ is defined as the radius of NSs and $M=M(R)$ is the gravitational mass. We have taken $G=c=1$ in Eq. (19). The TOV equation also gives us the mass density profile $\rho(r)$ of an NS; note that in the following text $\rho$ denotes mass density not $\rho$ meson field. Then the moment of inertia $I(r_{1},r_{2})$ can be calculated under rigid-body rotation,
@@
The total moment of inertia is $I(0,R)$. The superfluid moment of inertia is obtained by simply replacing $\rho$ with $(1-Y_{p})\rho$ in Eq. (20), where $Y_{p}$ is the proton fraction given by EoS.
Fig. 4 shows the mass-radius relations and mass profiles calculated from different EoSs. As shown in Fig. 4(a), all EoSs satisfy current observations of NS mass and radius. The maximum NS mass $M_{\rm{TOV}}$ corresponding to each EoS can be found in Table 3. The interaction in the isovector channel affects the radius of typical NSs ($\sim 1.4M_{\odot}$): the larger $L_{0}$, the larger the radius. In Fig. 4(b), we can see that the relations between central density of NSs $\rho_{\rm{c}}$ and NS masses are primarily determined by the interaction in the isoscalar channel for those NSs which are heavier than the typical NS.

### v2

@@ Abstract
We connect nuclear forces to one of the most notable irregular behaviors observed in pulsars, already detected in approximately 6% known pulsars, with increasingly accurate data expected from upcoming high-precision timing instruments on both ground and space. Built on Shang & Li (2021), we conduct a case study on the 2000 glitch of the Vela pulsar. For our purpose, we adopt the Relativistic Mean Field (RMF) model as the theoretical many-body framework to describe nuclear systems. We refit three representative RMF parameter sets (DD-ME2, PKDD, NL3), considering the uncertainties in nuclear matter saturation properties. Utilizing the resulting star structure, composition and nucleon properties in the medium obtained in a consistent manner, we calculate the pinning energy of superfluid vortex in the nuclear lattice in the inner crust. This leads to the evolution of associated pinning force that acts on the vortex, which can be confronted with observed glitch amplitude and short-time relaxation in the 2000 Vela glitch event, following the snowplow model of pulsar glitch. We discuss how the vortex configuration and pinning properties depend on the nuclear parameters, and find an interesting and dominant role of the nuclear symmetry energy slope on pinning strength.
@@ 1 Introduction
This paper is organized as follows. In Sec. 2.12.5, the theoretical framework for constructing unified EoS, the composition of NSs, the neutron pairing gap, and the pinning energy are given, respectively. Sec. 2.6 is dedicated to formulating the snowplow model. In Sec. 3, we demonstrate the dependence of pinning energy and pinning force on the effective interactions and polarization, and present the constraint on the nuclear parameters and Vela mass with the 2000 Vela glitch observation. Some necessary discussions are also given in Sec. 3. Finally, a brief summary is given in Sec. 4.
@@ 2.1 RMF framework for nuclear matter and refitting parameter sets
For describing nuclear matter, we consider the baryons interacting with each other through the exchange of isoscalar scalar and vector mesons ($\sigma$ and $\omega$), isovector vector meson ($\rho$) in the RMF model. The Lagrangian density that describes the systems with time-reversal symmetry can be written as (Boguta & Bodmer, 1977; Reinhard, 1989; Fuchs et al., 1995; Reinhard, 1989; Ring, 1996; Bender et al., 2003; Nikšić et al., 2011; Oertel et al., 2017): {widetext}
$\mathcal{L}=$ $\sum_{N=n,p}\bar{\psi}_{N}\left\{\gamma^{\mu}\left[i\partial_{\mu}-g_{\omega N}\omega_{\mu}-g_{\rho N}\boldsymbol{\rho}_{\mu}\boldsymbol{\tau}_{N}-q_{N}A_{\mu}\right]-\left[M_{N}-g_{\sigma N}\sigma\right]\right\}\psi_{N}$ (1)
@@
where $\boldsymbol{\tau}_{N}$ and $q_{N}$ are the Pauli matrices and charges of nucleons, and $M_{N}$ and $m_{l}$ represent the nucleon and lepton masses, respectively. $g_{mN}$ is the coupling constant between the meson $m$ and the nucleon $N$. $\psi_{N(l)}$ is the Dirac field of the nucleons or the leptons. $\sigma$, $\omega_{\mu}$, and $\boldsymbol{\rho}_{\mu}$ denote the quantum fields of mesons. The field tensors of $\omega$, $\rho$, and photon are
@@
The interaction term in the RMF Lagrangian Eq. (1) depends on the nucleon-meson coupling constants that are usually determined by fitting nuclei or nuclear-matter properties. For a quantitative description of nuclear matter and finite nuclei, one needs to include a medium dependence of the effective mean-field interactions accounting for higher-order many-body effects. The medium dependence can either be introduced by including non-linear (NL) meson self-interaction terms in the Lagrangian, or by assuming an explicit density dependence (DD) for the meson-nucleon couplings. In the NLRMF model, self-coupling terms and cross-coupling terms for the mesons are introduced in $U(\sigma,\omega_{\mu},\boldsymbol{\rho}_{\mu})$ with the flexibility of adjusting the density dependence of the symmetry energy, defined based on the parabolic approximation of the binding energy per nucleon $E/A=\varepsilon/\rho-M_{N}$: $E/A(\rho_{\rm B},\beta)=E/A(\rho_{\rm B},\beta=0)+E_{\rm sym}(\rho_{\rm B})\beta^{2}+...$ with $\beta=(\rho_{n}-\rho_{p})/\rho_{\rm B}$ denoting the isospin asymmetry. Density-dependent coupling parameters in the DDRMF model correspond to $U(\sigma,\omega_{\mu},\boldsymbol{\rho}_{\mu})$ in the NLRMF model. The density dependence of the coupling can be in phenomenological functional form, with parameters adjusted to experimental data. For example, following the Typel-Wolter ansatz (Typel & Wolter, 1999), the density dependence of the coupling parameters are
@@
The RMF model provides an excellent tool to study the properties of finite nuclei and infinite nuclear matter (Walecka, 1974; Glendenning, 1996; Shen, 2002; Dutra et al., 2014; Li et al., 2020; Rong et al., 2021; Tu & Zhou, 2022; Sun et al., 2023; Zhu et al., 2023). See Dutra et al. (2014) for a recent comparison of different RMF parameterizations. The extensions of several effective interactions, e.g., DD-ME2 (Lalazissis et al., 2005; Li & Sedrakian, 2019), PKDD (Long et al., 2004), and NL3 (Schaffner-Bielich & Gal, 2000; Wu et al., 2021), can be utilized to study the impact of nuclear symmetry energy on the pinning properties, with different density-slope of the symmetry energy at saturation density, $L_{0}=3\rho_{0}(\frac{\mathrm{d}E_{\rm sym}}{\mathrm{d}\rho_{\rm B}})|_{\rho_{0}}$. Except for pulsar glitch, the other NS phenomena, e.g., cooling (Newton et al., 2013) and asteroseismology (Neill et al., 2024), have been used to constrain $L_{0}$.
Table 1 lists the characteristic coefficients of nuclear matter for the original effective interactions DD-ME2, PKDD, and NL3. The original DD-ME2 is a widely-adopted interaction that effectively describes the properties of finite nuclei and nuclear matter, having the smallest symmetry energy slope $L_{0}=$ 51.26 MeV among the three effective interactions. In contrast, the original NL3 is characterized by an excessively large symmetry energy slope of $L_{0}=$ 118.5 MeV.
@@
Table 2 presents the coupling parameters $g_{\rho}(\rho_{0})$ and $a_{\rho}$ for different effective interactions with $L_{0}=$30, 40, 60, and 80 MeV. NL3 with $L_{0}=$ 30 MeV is excluded from our work due to its non-monotonic behavior in the EoS.
@@ 2.2 Unified description of stellar crust and core
For uniform nuclear matter in the core, due to the translational invariance of the system, all derivative terms in Eqs. (3)–(6) vanish. Using the Thomas-Fermi approximation, the nucleon Dirac field can be considered as plane wave, leading to a nucleon single-particle energy,
@@
The value of $E_{N}(k)$ at the Fermi surface is the corresponding chemical potential. $\Sigma_{R}$ is the rearrangement term that originated from the density dependence of the coupling constants between nucleons and mesons, $\Sigma_{R}=0$ for the NLRMF model. By combining the $\beta$-equilibrium condition, the charge neutrality condition, and the nucleon number conservation condition for uniform nuclear matter in the core, the nucleon Fermi momentum and meson fields at a given number density can be self-consistently solved.
Furthermore, the energy density $\varepsilon$ and pressure $P$ can be calculated by using the energy–momentum tensor,
@@
For the non-uniform matter of the outer and inner crusts, the microscopic nuclear structure can be described by the Wigner-Seitz (WS) cell with different geometric parameters. The energy density $\varepsilon_{\rm{non}}$ is calculated by integrating the whole WS cell,
@@
Apply the spherical WS approximation and reflective boundary condition, the distributions of meson fields and densities of nucleons and leptons are obtained by iteratively solving the nonlinear system of equations containing Eqs. (3-6), $\beta$-equilibrium, charge neutrality, and nucleon number conservation conditions in the spherical symmetry one-dimensional geometry. In each iteration, the meson field is expanded by fast cosine transformation to improve the computational efficiency. The non-uniform structure is stable after fulfilling the constancy of chemical potentials,
@@
The vertical lines on the left and right sides of Fig. 1 represent the outer-inner crust transition densities $\rho_{\rm{oi}}$ and crust-core transition densities $\rho_{\rm{cc}}$ derived from different effective interactions, respectively. Table 3 lists their exact values. The outer-inner crust transition density $\rho_{\rm{oi}}$ is fixed by the chemical potential of neutron exceeding its rest mass $M_{N}$, while the crust-core transition density $\rho_{\rm{cc}}$ is determined by $\varepsilon_{\rm{uni}}$ being less than $\varepsilon_{\rm{non}}$. It is clearly that as $L_{0}$ increases, $\rho_{\rm{oi}}$ increases while $\rho_{\rm{cc}}$ decreases, consistent with previous works (Fantina et al., 2016).
@@ 2.3 Global properties and composition of neutron stars
Figure 4: (a) $M$-$R$ relations and (b) mass profiles calculated with the unified EoSs of DD-ME2 (solid lines), PKDD (dashed lines), and NL3 (dotted lines) with the different symmetry energy slope $L_{0}=$30, 40, 60, and 80 MeV. The mass-radius measurements from GW observations for GW170817 (Abbott et al., 2017) and NICER missions for PSR J0030+0451 (Vinciguerra et al., 2024) and PSR J0740+6620 (Riley et al., 2021) are shaded in panel (a). The dashed horizontal line presents $2M_{\odot}$ in panel (b).
Now we discuss the dependence of droplet size $R_{\rm{d}}$ and WS cell size $R_{\rm{WS}}$ on effective interactions, particularly $L_{0}$. Fig. 3 illustrates $R_{\rm{d}}$ and $R_{\rm{WS}}$ calculated using DD-ME2 with $L_{0}=$30, 60, and 80 MeV as a function of $\rho_{\mathrm{B}}$. At lower densities, $L_{0}$ has a minor effect on $R_{\rm{d}}$ and $R_{\rm{WS}}$. However, near the crust-core transition densities, we can observe a dramatic increase in both $R_{\rm{d}}$ and $R_{\rm{WS}}$; this effect becomes more pronounced with smaller $L_{0}$ and tends to disappear at $L_{0}=$ 80 MeV. The effects of symmetry energy on $R_{\rm{d}}$ and $R_{\rm{WS}}$ can be explained by the surface tension. Oyamatsu & Iida (2007) has showed that the surface tension decreases with increasing $L_{0}$. Consequently, droplets calculated at lower $L_{0}$ can maintain inhomogeneous structures under higher density and stronger Coulomb force due to the larger surface tension, thus further supporting a higher transition density and a larger droplet size. The upper panel of Fig. 3 for $L_{0}=$ 30 MeV suggests the possible existence of the giant droplet and extremely dense inhomogeneous structure near the crust-core transition density, which could significantly reduce pinning force.
Both the global properties and composition of NSs are essential to simulations of pulsar glitches. Using the unified EoS shown in Fig. 1 as basic input, the global properties of NSs, e.g., mass-radius ($M$-$R$) relation and mass profile, can be obtained by solving the Tolman–Oppenheimer–Volkoff (TOV) equation (Tolman, 1939; Oppenheimer & Volkoff, 1939),
@@
where $r$ is the distance from the center of NSs. Given a central density $\rho_{c}$, the TOV equation is integrated from $r=0$ to $r=R$ where the pressure is zero. $R$ is defined as the radius of NSs and $M=M(R)$ is the gravitational mass. We have taken $G=c=1$ in Eq. (19). The TOV equation also gives us the mass density profile $\rho(r)$ of an NS; note that in the following text $\rho$ denotes mass density not $\rho$ meson field. Then the moment of inertia $I(r_{1},r_{2})$ can be calculated under rigid-body rotation,
@@
The total moment of inertia is $I(0,R)$. The superfluid moment of inertia is obtained by simply replacing $\rho$ with $(1-Y_{p})\rho$ in Eq. (20), where $Y_{p}$ is the proton fraction for a given EoS.
Fig. 4 shows the $M$-$R$ relations and mass profiles calculated from different EoSs. As shown in Fig. 4(a), all EoSs satisfy current observations of NS mass and radius. The maximum NS mass $M_{\rm{TOV}}$ corresponding to each EoS can be found in Table 3. The interaction in the isovector channel affects the radius of typical NSs ($\sim 1.4M_{\odot}$): the larger $L_{0}$, the larger the radius. In Fig. 4(b), we can see that the relations between central density of NSs $\rho_{\rm{c}}$ and NS masses are primarily determined by the interaction in the isoscalar channel for those NSs which are heavier than the typical NS.